• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 04:41:22 on June17,2002

    Entry number 86223


    keyword=Bogie problems & Q1-L trip

    Problems when we wanted to change to kinematics Q2_f10r

    1. First, we accidentally change L-arm angle. We realized the mistake before the
    spectrometer moved 0.25 degree, then put it back to floor mark 21, vernier -83.5
    which was exactly where it had been (to the extent the system reproduces). Sorry !

    2. Second problem, we could not enable the R-arm bogies. We tried to reboot the
    bogie VME control crate. Didn't help. We paged Scott Spiegel and he suggested
    we power cycle the bogies and a UPS near the center bogie. Took access to do
    that and it worked fine.

    3. Meanwhile Q1 on Left HRS tripped. We reset it.

    Back in business for Q2_f10r